# JAX-MPM: A Learning-Augmented Differentiable Meshfree Framework for GPU-Accelerated Lagrangian Simulation and Geophysical Inverse Modeling

### v1

@@ 6 Conclusion
We presented JAX-MPM, a fully differentiable, GPU-accelerated Material Point Method (MPM) framework tailored for large-deformation simulation and inverse modeling in geomechanics and solid mechanics. Built entirely on the JAX ecosystem, the framework leverages modern differentiable programming capabilities, including automatic differentiation, just-in-time (JIT) compilation, and vectorized GPU execution, to support scalable, end-to-end optimization pipelines. JAX-MPM integrates naturally with neural network libraries, enabling hybrid physical–ML workflows for data-driven modeling and field reconstruction.
The framework employs a unified inverse modeling approach that accommodates both particle-based and region-based observations, allowing flexible inference of physical parameters from diverse data sources. We validated JAX-MPM on a suite of forward benchmarks, including dam-break problems and granular collapses in both 2D and 3D, demonstrating its accuracy, efficiency, and scalability. Compared to traditional CPU-based implementations in C++ or Fortran, JAX-MPM achieves significant speedups through GPU acceleration, while providing native support for gradient-based optimization and machine learning integration. To evaluate inverse modeling performance, we conducted several tasks such as initial condition reconstruction and spatially varying friction estimation These results position JAX-MPM as a robust and extensible platform for differentiable simulation in geomechanics.
Looking ahead, several opportunities remain for further development. First, while checkpointing helps mitigate memory usage during training, the backward pass in high-resolution simulations remains demanding—particularly with small time steps or dense grids. Enhancing memory management and gradient computation efficiency will be key to scaling up optimization. Second, modeling non-smooth or highly nonlinear materials—such as Drucker–Prager plasticity—can cause instability or undefined gradients during automatic differentiation. Addressing this may require customized backward kernels or hybrid constitutive models that blend physics-based laws with neural approximators. Lastly, extending the framework to support implicit time integration could improve numerical stability and enable longer simulation horizons. While feed-forward neural networks are currently used in JAX-MPM to approximate unknown parameter fields, the framework’s general differentiability also allows for alternative approaches such as symbolic regression or hybrid formulations (e.g., NIM [33], enrichment functions incorporating prior knowledge).
@@
Overall, JAX-MPM advances differentiable physics in continuum mechanics by bridging high-fidelity simulation with scalable gradient-based optimization. It offers a flexible and efficient foundation for future research in inverse modeling, hybrid physical–ML systems, and learning-enabled digital twins for geohazard analysis.
@@ Appendix A Stress update for Drucker–Prager elastoplasticity
This appendix details the implementation of Drucker–Prager constitutive model. The model parameters are defined as
@@
The constitutive model follows a rate-independent, return-mapping algorithm with Drucker–Prager-type shear and tensile yielding. The Jaumann stress rate ensures objectivity under large deformation. The procedure is as follows:
@@
    Elastic trial stress. For clarity, we reformulate the stress tensor in Eq. (21) using index notation. The trial stress at $t^{n+1}$ is evaluated by combining the rotated stress and the deformation increment:
    ${}^{*}\sigma_{ij}^{n+1}=\sigma_{ij}^{Rn}+C_{ijkl}\Delta d_{kl}$ (37)
    where $C_{ijkl}$ is the tangential stiffness tensor of the material, $\sigma_{ij}^{Rn}$ accounts for material rotation via the Jaumann rate:
    $\sigma_{ij}^{Rn}=\sigma_{ij}^{n}+\sigma_{ik}^{n}\Delta\omega_{jk}+\sigma_{jk}^{n}\Delta\omega_{ik}$ (38)
@@
    Compute trial invariants. The spherical (mean) and effective shear stresses of the trial stress are given by:
    ${}^{*}\sigma_{m}^{n+1}$ $=\frac{1}{3}{}^{*}\sigma_{kk}^{n+1}$ (39)
    ${}^{*}\tau^{n+1}$ $=\sqrt{\frac{1}{2}{}^{*}s_{ij}^{n+1}{}^{*}s_{ij}^{n+1}}$ (40)
    where ${}^{*}s_{ij}={}^{*}\sigma_{ij}^{n+1}-{}^{*}\sigma_{m}^{n+1}$ is the deviatoric components of trial stress.
@@
    Evaluate yield conditions. Plastic yielding is governed by three criteria:
    $f^{s}$ $={}^{*}\tau^{n+1}-k_{\varphi}+q_{\varphi}{}^{*}\sigma_{m}^{n+1}$ (41)
    $f^{t}$ $={}^{*}\sigma_{m}^{n+1}-\sigma^{t}$ (42)
    $h$ $={}^{*}\tau^{n+1}-\tau^{P}-\alpha^{P}({}^{*}\sigma_{m}^{n+1}-\sigma^{t})$ (43)
    where $f^{s}$ is the Drucker–Prager shear yield function, and $f^{t}$ represents a tensile cutoff. And $h$ [102] is a auxiliary function with the constants $\tau^{P}$ and $\alpha^{P}$ given by:
    $\tau^{P}=k_{\varphi}-q_{\varphi}\sigma^{t}\quad\text{and}\quad\alpha^{P}=\sqrt{1+q_{\varphi}^{2}}-q_{\varphi}$ (44)
    The trial stress state is classified into the following three zones:
@@
    In Zone II, the stress is returned to the shear yield surface $f^{s}=0$ by a radial return rule, assuming incompressible plastic flow. The updated stress components are:
    $\sigma_{m}^{n+1}$ $={}^{*}\sigma_{m}^{n+1}-Kq_{\psi}\Delta\lambda^{s}$ (48)
    $\tau^{n+1}$ $=k_{\varphi}-q_{\varphi}\sigma_{m}^{n+1}$ (49)
    $\sigma_{ij}^{n+1}$ $=\frac{\tau^{n+1}}{{}^{*}\tau^{n+1}}\,{}^{*}s_{ij}^{n+1}+\sigma_{m}^{n+1}\delta_{ij}$ (50)
@@
    $\Delta\varepsilon_{eq,p}=\Delta\lambda^{s}\sqrt{\frac{1}{3}+\frac{2}{9}q_{\psi}^{2}}$ (51)
    The proportionality constant $\Delta\lambda^{s}$ is obtained by enforcing consistency with the yield surface:
    $\Delta\lambda^{s}=\frac{f^{s}}{G+Kq_{\varphi}q_{\psi}}$ (52)
@@
    In Zone III, tensile failure occurs when the trial mean stress exceeds $\sigma^{t}$ and the state lies outside the transition surface. The material can no longer sustain tension, and the stress is corrected to cap the mean stress at $\sigma^{t}$.
    The corrected stress and equivalent plastic strain are computed as:
    $\sigma_{ij}^{n+1}$ $={}^{*}\sigma_{ij}^{n+1}+\left(\sigma^{t}-{}^{*}\sigma_{m}^{n+1}\right)\delta_{ij}$ (53)
    $\Delta\varepsilon_{eq,p}$ $=\frac{\sqrt{2}}{3}\Delta\lambda^{t}$ (54)
    The proportionality constant $\Delta\lambda^{t}$ is derived from the consistency condition $f^{t}=0$:
    $\Delta\lambda^{t}=\frac{{}^{*}\sigma_{m}^{n+1}-\sigma^{t}}{K}$ (55)
Appendix B Velocity transfer scheme
JAX-MPM supports multiple grid-to-particle (G2P) transfer schemes, including PIC, FLIP, APIC, and TPIC. Among them, APIC [55] and TPIC [71] retain the same velocity update formulation as PIC for computing the particle velocity $\bm{v}_{p}^{n}$, but additionally incorporate particle-level information such as affine velocity fields or velocity gradients to reduce numerical dissipation and enhance stability. These additional particle quantities—such as the affine matrix $\bm{B}_{p}^{n}$ in APIC or the velocity gradient $\nabla\bm{v}_{p}^{n}$ in TPIC—are evaluated during the G2P step and reused in the subsequent P2G step to compute the nodal velocity field.
@@ Appendix B Velocity transfer scheme
In TPIC, the affine contribution is replaced by a particle velocity gradient $\nabla\bm{v}_{p}^{n}$, which is derived from prior particle and grid states. The nodal velocity is then computed as:

### v2

@@ 6 Conclusion
This study presents JAX-MPM, a fully differentiable, GPU-accelerated material point method (MPM) framework tailored for large-deformation predictive simulation and inverse modeling in continuum mechanics, with a focus on geophysical flow applications. Built entirely on the JAX ecosystem, the framework leverages modern differentiable programming capabilities, including automatic differentiation, JIT compilation, and vectorized GPU execution, to support scalable, end-to-end optimization pipelines. JAX-MPM also integrates naturally with neural network libraries, enabling hybrid physical–ML workflows for data-driven modeling and field reconstruction.
We validated JAX-MPM on a suite of forward benchmarks, including 2D and 3D dam-break problems and granular collapses, demonstrating high accuracy, computational efficiency, and scalability. Compared to conventional CPU-based implementations in C++ or Fortran, JAX-MPM achieves substantial speedups (up to 100×) by exploiting GPU acceleration. Using JAX-MPM, we further investigated the influence of initial geometry (e.g., aspect ratio) and material properties (e.g., internal friction angle) on granular flow dynamics. The results show strong agreement with theoretical predictions and experimental observations, underscoring the effectiveness of the framework for realistic geophysical flow simulation and its potential to advance the understanding and design of processes.
Beyond forward simulation, JAX-MPM enables inverse modeling through a unified formulation that supports both particle-based (Lagrangian) and region-based (Eulerian) supervision. To evaluate this capability, we conducted several inverse analysis tasks, including: (1) reconstruction of initial velocity fields, both parametric and non-parametric (neural network-based), from sparse trajectory or velocity monitor data; (2) estimation of spatially varying friction coefficients along the base of a dam-break domain, even under contact-dominated dynamics. These experiments demonstrate that JAX-MPM can recover localized and spatially varying parameters from limited observations with high fidelity. To the best of our knowledge, this represents the first demonstration of a differentiable physics-based solver applied to large-scale Lagrangian geophysical flows for inverse frictional parameter estimation. This makes it a versatile tool for real-world geomechanical applications, where direct measurement is often infeasible and observational data is sparse.
Furthermore, we showcase the integration of feed-forward neural networks within JAX-MPM to approximate unknown parameter fields (Section 5.2). The successful reconstruction illustrates the potential of neural parameterizations for capturing complex functional relationships, benefiting from universal approximation and smooth differentiability. The framework’s general differentiability also opens opportunities for incorporating alternative modeling paradigms such as symbolic regression, Gaussian process regression, or hybrid symbolic–neural formulations (e.g., NIM [32]) when prior knowledge is available.
@@
In summary, JAX-MPM advances the field of differentiable physics for continuum mechanics by bridging high-fidelity simulation with scalable gradient-based optimization. It offers a flexible and efficient foundation for future research in inverse modeling, learning-enabled simulation, and digital twins for geohazard analysis, natural hazard mitigation, and intelligent design of geotechnical systems.
@@ Appendix A Stress update for Drucker–Prager elastoplasticity
This appendix details the Drucker–Prager constitutive model and its rate-independent return mapping update with Jaumann objective rate. The model parameters are defined as
@@
We adopt a rate-independent return-mapping algorithm that incorporates Drucker–Prager shear yielding and a tensile failure cutoff, with objectivity ensured via the Jaumann stress rate. The procedure involves the following steps [51]:
@@
    Elastic trial stress. For clarity, we reformulate the stress tensor in Eq. (21) and other quantities using index notation. The trial stress at $t^{n+1}$ is evaluated by combining the rotated stress and the deformation increment:
    $^{*}\sigma_{ij}^{n+1}=\sigma_{ij}^{Rn}+C_{ijkl}\Delta d_{kl},$ (41)
    where $C_{ijkl}$ is the tangential stiffness tensor of the material, $\sigma_{ij}^{Rn}$ is the Jaumann-rotated stress from the previous step:
    $\sigma_{ij}^{Rn}=\sigma_{ij}^{n}+\sigma_{ik}^{n}\Delta\omega_{jk}+\sigma_{jk}^{n}\Delta\omega_{ik}.$ (42)
@@
    Trial invariants. The trial spherical (mean) and effective shear stresses of the trial stress are given by:
    ${}^{*}\sigma_{m}^{n+1}$ $=\frac{1}{3}{}^{*}\sigma_{kk}^{n+1},$ (43)
    ${}^{*}\tau^{n+1}$ $=\sqrt{\frac{1}{2}{}^{*}s_{ij}^{n+1}{}^{*}s_{ij}^{n+1}},$ (44)
    where $^{*}s_{ij}^{n+1}={}^{*}\sigma_{ij}^{n+1}-{}^{*}\sigma_{m}^{n+1}$ is the deviatoric components of trial stress.
@@
    Yield conditions. Plastic yielding is governed by the following three functions:
    $\text{Shear failure yield function:}\quad f^{s}$ $={}^{*}\tau^{n+1}-k_{\varphi}+q_{\varphi}{}^{*}\sigma_{m}^{n+1}$ (45)
    $\text{Tension failure yield function:}\quad f^{t}$ $={}^{*}\sigma_{m}^{n+1}-\sigma^{t}$ (46)
    $\text{Auxiliary transition function:}\quad h$ $={}^{*}\tau^{n+1}-\tau^{P}-\alpha^{P}({}^{*}\sigma_{m}^{n+1}-\sigma^{t})$ (47)
    where the auxiliary function $h$ [101] defines a transition boundary between shear and tensile failure domains, with constants $\tau^{P}$ and $\alpha^{P}$ given by:
    $\tau^{P}=k_{\varphi}-q_{\varphi}\sigma^{t}\quad\text{and}\quad\alpha^{P}=\sqrt{1+q_{\varphi}^{2}}-q_{\varphi}.$ (48)
    Based on the evaluated yield functions, the trial stress state is classified into three zones:
@@
    In Zone II, the trial stress violates the shear yield condition $f^{s}>0$, indicating plastic shear deformation. The stress is projected back to the shear yield surface $f^{s}=0$ using a radial return mapping algorithm under the assumption of incompressible plastic flow. The updated stress components are:
    $\sigma_{m}^{n+1}$ $={}^{*}\sigma_{m}^{n+1}-Kq_{\psi}\Delta\lambda^{s},$ (52)
    $\tau^{n+1}$ $=k_{\varphi}-q_{\varphi}\sigma_{m}^{n+1},$ (53)
    $\sigma_{ij}^{n+1}$ $=\frac{\tau^{n+1}}{{}^{*}\tau^{n+1}}\,{}^{*}s_{ij}^{n+1}+\sigma_{m}^{n+1}\delta_{ij},$ (54)
@@
    $\Delta\varepsilon_{eq,p}=\Delta\lambda^{s}\sqrt{\frac{1}{3}+\frac{2}{9}q_{\psi}^{2}}.$ (55)
    The plastic multiplier $\Delta\lambda^{s}$ is determined by the consistency condition: $f^{s}=0$
    $\Delta\lambda^{s}=\frac{f^{s}}{G+Kq_{\varphi}q_{\psi}}.$ (56)
@@
    In Zone III, where when the trial mean stress exceeds $\sigma^{t}$ and the state lies outside the transition surface, the tensile failure occurs. A plastic correction is applied to cap the mean stress at $\sigma^{t}$. The updated stress and equivalent plastic strain are:
    $\sigma_{ij}^{n+1}$ $={}^{*}\sigma_{ij}^{n+1}+\left(\sigma^{t}-{}^{*}\sigma_{m}^{n+1}\right)\delta_{ij},$ (57)
    $\Delta\varepsilon_{eq,p}$ $=\frac{\sqrt{2}}{3}\Delta\lambda^{t},$ (58)
    where the plastic multiplier $\Delta\lambda^{t}$ is derived from the consistency condition $f^{t}=0$:
    $\Delta\lambda^{t}=\frac{{}^{*}\sigma_{m}^{n+1}-\sigma^{t}}{K}.$ (59)
Appendix B Velocity transfer schemes
JAX-MPM supports multiple grid-to-particle (G2P) transfer schemes, including PIC, FLIP, APIC, and TPIC. Among them, APIC [54] and TPIC [70] retain the same velocity update formulation as in PIC for computing the particle velocity $\bm{v}_{p}^{n}$, but additionally incorporate particle-level information such as affine velocity fields or velocity gradients to reduce numerical dissipation and enhance stability. These additional particle quantities (such as the affine matrix $\bm{B}_{p}^{n}$ in APIC or the velocity gradient $\nabla\bm{v}_{p}^{n}$ in TPIC) are evaluated during the G2P step and then used in the subsequent P2G step to compute the nodal velocity field.
@@ Appendix B Velocity transfer schemes
In TPIC, the affine contribution is replaced by the particle velocity gradient $\nabla\bm{v}_{p}^{n}$, which is derived from prior particle and grid states. Thus, the nodal velocity is computed as: